You are given an RSA crytosystem with primes of p=101, q=701 and a private exponent of d=23, what is the complete public key information that needs to be published? (20 marks)
1
Expert's answer
2017-07-27T15:49:06-0400
The values of p=101 and q=701 help calculate N=p*q=70801, and also a number r=(p-1)(q-1)= 70000, which is very important. Now to find two numbers e and d whose product is a number equal to 1 mod r: de mod r=23, so, find e. e=d-1 mod r=6087
So, published must be: e=6087; N=70801.